2. (10 pts.) Show that either g(n) Ofn)) or frn) -O(g(n) for the following. g(n)n2 + 7 n, f(n)-ns_2㎡ a. b. g(n)= 2n + 4,f(n)=61g(n*)
media%2Fb20%2Fb208f7ae-7603-4472-9237-d8
Answer #1

Solution:

before starting let's have a look at Big-O first.

u) n) s upper bound in) 2 3 4

2)

a)

we\ have\ g(n) = n^2 + 7n, and\ f(n)= n^3-2n^2

comparing f(n) and g(n) we can see that:

n^2 + 7n <= c*n^3-2n^2

for c>0, the above inequality holds true

so, g(n)= O(f(n))

b)

we\ have\ g(n) = 2n + 4, and\ f(n)= 6\ lg(n^2)

after applying logarithmic property we can wrtie f(n) as

f(n)= 12 lg n

comparing f(n) and g(n),

we can see that

f(n)<= c*g(n)

for c> 0

This means that g(n)= O(f(n))

I hope this helps if you find any problem. Please comment below. Don't forget to give a thumbs up if you liked it. :)